Cinco De Mayo has a target debt-equity ratio of 0.56. (Hint: if you're not sure what to do about "debt-equity ratio", watch my YouTube video on that topic!) The yield to maturity on its bonds is 11 percent. Its cost of equity is 20 percent. The corporate income tax rate is 32 percent. Calculate the WACC for this company.

Multiple Choice

  • 11.97%

  • 12.99%

  • 15.51%

  • 16.28%

  • 14.73%
Answer #1

WACC = weight of debt * after tax cost of debt + weight of equity * cost of equity

debt/equity = 0.56

and debt + equity = 1

=>

equity = 25/39

debt = 14/39

WACC = 14/39 * 11% * (1-0.32) + 25/39 * 20%

= 15.51%